P=100-0.5Q where Q=q1+q2, TC1=5q1,TC2=0.5q2 find betrand equilibrum
1
Expert's answer
2019-02-04T09:37:08-0500

5q1=0.5q2

q_2=(5q_1)/0.5=10q_1

Q=q1+q2=11q1

Equilibrium price will be following:

P=100-0.5×11q1=100-5.5q1=100-0.55q2